OpenStudy (anonymous):

I am trying to find the nth of a sequence : a. 1, 3, 6, 10, 15 ..i need to know how to do this

OpenStudy (anonymous):

what value of n are we trying to find? in other words, which term are you trying to find?

OpenStudy (anonymous):

looks like you are adding one more each time. or you can figure out an equation

OpenStudy (anonymous):

1+2=3 3+3=6 6+4=10 10+5=15 etc so next one would be 15+6=21 21+7=28 etc

OpenStudy (anonymous):

This is the question..What is the nth term in each of the numerical patterns.and those are the numbers

OpenStudy (anonymous):

I don't need to know what comes next, but a formula

OpenStudy (anonymous):

\[a_{n}=a_{n-1}+n\]

OpenStudy (anonymous):

johnnyrocket has it but there is a closed form as well i will see if i can find it

OpenStudy (anonymous):

If we are looking for a general formula for the sequence it would be \[a _{n}=a _{n-1}+n\] for all n>1 where\[n _{1}=1\]

OpenStudy (anonymous):

i have three differnet sequences so i need to learn how to plug in the numbers or me uwith a fomrula

OpenStudy (anonymous):

i think we want the closed form of this thing.

OpenStudy (anonymous):

i am getting there i think. give me a few

OpenStudy (anonymous):

((n+1)n)/2

OpenStudy (anonymous):

(n^2+n)/2

OpenStudy (anonymous):

i believe you but why the heck to i keep getting \[\frac{n^2-n}{2}\] which i know is wrong!

OpenStudy (anonymous):

maybe bad algebra? or bad method.

OpenStudy (anonymous):

once you have it it is clear. for some reason i kept getting minus sign can you show me how to get it right?

OpenStudy (anonymous):

How do you get a negative

OpenStudy (anonymous):

oh now i see it. i was summing \[\sum_{k=1}^nk=\frac{n(n+1)}{2}\] when i should have been summing \[\sum_{k=1}^{n-1}k=\frac{n(n-1)}{2}\]

OpenStudy (anonymous):

then \[n^2-\frac{n(n-1)}{2}\]will work. anyway that is what i did. was there a snappier way?

OpenStudy (anonymous):

my way was probably the donkey way. i am sure there is a snap way to do this, i forget how

OpenStudy (anonymous):

believe it or not i just saw this a couple years ago and now totally forget. solve these and recursions like \[a_n=a_{n-1}+2a_{n-2}\] etc

OpenStudy (anonymous):

i would love to see what you did to get this